how do I solve by addition. the second equation of the system has me stumped. How do I elimated one?
5x-7y=2
3x=4y

Answer by algebrahouse.com(1659)

"how do I solve by addition. the second equation of the system has me stumped. How do I elimate one?"
5x - 7y = 2
3x = 4y-------->3x - 4y = 0 {subtracted 4y from both sides in 2nd equation}

5x - 7y = 2-------> -15x + 21y = -6 {multiplied top equation by -3}
3x - 4y = 0-------> 15x - 20y = 0 {multiplied bottom equation by 5}

-15x + 21y = -6
15x - 20y = 0
--------------
y = -6 {added two new equations together and eliminated x}

5x - 7y = 2 {top original equation}
5x - 7(-6) = 2 {substituted -6, in for y, into top original equation}
5x + 42 = 2 {multiplied -7 by -6}
5x = -40 {subtracted 42 from both sides}
x = -8 {divided both sides by 5}

x = -8 and y = -6
www.algebrahouse.com